### v1

# Simulation-based Approach for Fast Optimal Control of a Stefan Problem with Application to Cell Therapy

Abstract
This article describes a new, efficient way of finding control and state trajectories in optimal control problems by transformation into a system of differential-algebraic equations (DAEs). The optimal control and state vectors can be obtained via simulation of the resulting DAE system with the selected DAE solver, eliminating the need for an optimization solver. Our simulation-based framework was demonstrated and benchmarked against various optimization-based approaches via four case studies associated with optimization and control of a Stefan problem with application to cell therapy. The simulation-based approach is faster than every optimization-based method by more than an order of magnitude while giving the same level of accuracy in all cases. The proposed technique offers an efficient and reliable framework for optimal control, serving as a promising alternative to the traditional techniques in applications where speed is crucial, e.g., real-time online model predictive control.
keywords:
[…]
thanks: Corresponding author: Richard D. Braatzthanks: This manuscript is an extension of a paper published in the Proceedings of the American Control Conference (Srisuma et al., 2024).
[…]
1 Introduction
A Stefan problem describes the evolution of a moving interface during phase change, e.g., freezing and melting (Carslaw and Jaeger, 1959; Bird et al., 2002). Different Stefan problem formulations have been applied to study various industrial and natural systems, including polymorphous materials (Tao, 1979), steel casting (Hill and Wu, 1994; Petrus et al., 2010), biological tissue (Rabin and Shitzer, 1995, 1997), alloy formation (Brosa Planella et al., 2019, 2021), glaciation (Mikova et al., 2017), phase change materials (Brezina et al., 2018), cryopreservation (Dalwadi et al., 2020), and cell therapy (Srisuma et al., 2023b). Numerical techniques have been developed for implementing and simulating the Stefan problems (Meyer, 1971; Velardi and Barresi, 2008; Kurbatova and Ermolaeva, 2019; Gusev et al., 2021; Srisuma et al., 2023b).
Optimal control of Stefan problems was extensively investigated and proven useful for various industrial applications over the past few decades, with many different objective functions, constraints, and controls (manipulated variables) considered. Some examples are the control of the heating process to satisfy the heating speed and thermoelastic stress constraints by manipulating the furnace temperature (Roubíček, 1986), the stabilization of the moving boundary and temperature/concentration fields by varying the heat flux (Pawlow, 1987), the maximization of the amount of melted solid in the melting process via controlling the heat flux (Silva Neto and White, 1994), the control of the moving boundary to follow the desired path in the solidification process by manipulating the wall temperature (Hinze and Ziegenbalg, 2007), the control of the water level in the drainage basin by varying the discharge velocity (Miyaoka and Kawahara, 2008), the optimization of a thin-film drying process via manipulating the air temperature (Mesbah et al., 2014), and the minimization of the metallurgical length (ML) deviation in steel casting via controlling the boundary heat flux (Chen et al., 2019).
All the aforementioned studies rely on optimization algorithms/solvers to obtain the optimal control trajectories. The widely used approach is to replace the time-varying control vector by a finite-dimensional parameterization (e.g., a spline) and carry out numerical discretization to transform the dynamic optimization into a nonlinear algebraic program (Kishida et al., 2013; Scott et al., 2018; Nolasco et al., 2021, for example). This overall approach has many variations, including methods that sequentially switch between solving the numerical discretization of the underlying partial differential-algebraic equations (PDAEs) and running an algebraic optimizer, or simultaneously solving a single large sparse optimization with the numerical discretization equations as explicit constraints (Neuman and Sen, 1974; Tsang et al., 1975; Mellefont and Sargent, 1978; Sargent and Sullivan, 1978; Biegler, 2007; Nolasco et al., 2021). Numerous optimal control algorithms have been developed for efficient solutions to facilitate real-time applications such as model predictive control (e.g., see detailed discussions by Rodrigues et al. (2014) and Nolasco et al. (2021)). Alternatively, Berliner et al. (2022) showed that it is possible to reformulate some optimal control problems as a mixed continuous-discrete system of index-1 PDAEs. The PDAE system is then solved numerically by feeding the differential-algebraic equations (DAEs) obtained by spatial discretization into a DAE solver. In this approach, the optimal control vector is obtained directly by a DAE solver, without using any optimization solver, resulting in a highly computationally efficient solution to the optimal control problem. Recently, this approach was used to solve optimal control problems associated with a Stefan problem (Srisuma et al., 2024).
Although optimal control with Stefan problems has been explored for a wide range of processes, applications to cell therapy have not been available. Previous studies showed that accurate prediction, control, and optimization of cryopreservation and cell thawing can improve the viability and quality of the resulting cells, which directly benefits cell therapy (Seki and Mazur, 2008; Jang et al., 2017; Baboo et al., 2019; Hunt, 2019; Bojic et al., 2021; Cottle et al., 2022; Uhrig et al., 2022). These benefits and case studies therefore motivate the development of an efficient Stefan problem-based optimal control algorithm for cell therapy.
This article describes an optimal control algorithm via DAE reformulation and simulation, referred to as the simulation-based approach, for cell therapy applications. The main contributions of this work are to
  1. 1.
    derive the simulation-based approach for optimal control of a Stefan problem for the cell thawing process in cell therapy;
  2. 2.
    generalize the simulation-based approach for both index-1 and high-index DAE systems;
  3. 3.
    demonstrate and benchmark the simulation-based algorithm against various optimization-based optimal control algorithms; and
  4. 4.
    apply the proposed approach to efficiently and accurately solve several practical problems associated with optimization and control of cell thawing.
This article is organized as follows. Section 2 describes the cell thawing system and summarizes the mechanistic model and equations. Section 3 discusses various techniques for solving optimal control problems and introduces the simulation-based approach. Section 4 demonstrates the simulation-based approach via case studies. Finally, Section 5 summarizes the study.
2 Stefan Problem and Cell Thawing
The system used for the demonstration of our optimal control technique in this article concerns cell thawing, which is a process used in cell therapy before cells are introduced to the patients (Fig.  1). During thawing, energy is continuously supplied by a heater to thaw the material in a vial. A mechanistic model of cell thawing can be formulated as a Stefan problem (Srisuma et al., 2023b). Heat transfer in the solid and liquid domains can be described by the energy balance equation,
$\frac{1}{\alpha}\frac{\partial T}{\partial t}=\frac{\partial^{2}T}{\partial r^{2}}+\frac{1}{r}\frac{\partial T}{\partial r},$ (1)
where $T$ is the temperature, $r$ is the radial direction, $t$ is time, and $\alpha$ is the thermal diffusivity. Heat transfer associated with thawing at the moving solid-liquid interface is governed by the Stefan conditions
$\rho\Delta H_{f}\dfrac{ds}{dt}=k_{1}\dfrac{\partial T_{1}}{\partial r}-k_{2}\dfrac{\partial T_{2}}{\partial r},$ (2)
$T_{1}=T_{2}=T_{m},$ (3)
where $s$ is the interface position, $\Delta H_{f}$ is the latent heat of fusion, $T_{m}$ is the melting point, $\rho$ is the density, $k$ is the thermal conductivity, and the subscripts 1 and 2 denote the solid and liquid phases, respectively. The above equations are nondimensionalized and discretized using the finite difference scheme with the method of lines, with appropriate boundary conditions, resulting in
$\dfrac{d\Theta_{1}}{d\tau}=f_{1}(\Theta_{1},\Theta_{2},S),$ (4)
$\dfrac{d\Theta_{2}}{d\tau}=f_{2}(\Theta_{1},\Theta_{2},\Theta_{b},S),$ (5)
$\dfrac{dS}{d\tau}=f_{3}(\Theta_{1},\Theta_{2},S),$ (6)
$(\Theta_{1})_{n}=(\Theta_{2})_{0}=0,$ (7)
with the initial conditions
$\Theta_{1}(\tau_{0})=\Theta_{2}(\tau_{0})=0,$ (8)
$S(\tau_{0})=1,$ (9)
where $\Theta_{1}\in\mathbb{R}^{n}$ collects the solid temperature $(\Theta_{1})_{i}$ for $i=0,\dots,n-1$; $\Theta_{2}\in\mathbb{R}^{n}$ collects the liquid temperature $(\Theta_{2})_{j}$ for $j=1,\dots,n$; $S\in\mathbb{R}^{1}$ is the interface position; $f_{1}\in\mathbb{R}^{n}$, $f_{2}\in\mathbb{R}^{n}$, $f_{3}\in\mathbb{R}^{1}$ are the nonlinear functions; $n$ is the number of grid points in each domain (set to 20); $\tau\in[\tau_{0},\tau_{f}]$ is the dimensionless time; $\tau_{0}$ is the initial time; and $\tau_{f}$ is the final time. Here (4) and (5) are derived from (1), while (6) and (7) correspond to (2) and (3), respectively. We refer to Srisuma et al. (2023b) for the detailed derivation of all equations, parameter values, and model validation.
Figure 1: (A) In cell thawing, the vial containing biological cells frozen in ice is heated by the heater. (B) A schematic diagram showing the one-dimensional Stefan problem in a cylindrical coordinate system. The moving interface position is $S$. The solid temperature is $\Theta_{1}$. The liquid temperature is $\Theta_{2}$. The heater temperature is $\Theta_{b}$. All variables are written in dimensionless form.
In dimensionless form, the model consists of three main parts which describe the evolution of the solid temperature (4), the liquid temperature (5), and the interface position (6), respectively. The energy is continuously supplied to the system by the heater with the temperature $\Theta_{b}$. The system initially consists of pure solid, which corresponds to $S=1$; i.e., the interface position is at the outer boundary. As the thawing progresses, the interface position decreases and eventually becomes 0, indicating complete melting or pure liquid. The maximum temperature (dimensionless) is equal to 1, which is equivalent to $37\,^{\circ}$, the maximum temperature recommended for cell thawing (Baboo et al., 2019; Uhrig et al., 2022). The minimum temperature (dimensionless) is equal to 0, corresponding to the melting point of $-2\,^{\circ}$.
Finally, since the temperature varies spatially, we define the average temperature $\Theta_{\textrm{avg}}$ as
$\Theta_{\textrm{avg}}=\sum_{j=1}^{n}\dfrac{1}{n}{(\Theta_{2})_{j}}.$ (10)
In this case, only the liquid temperature $\Theta_{2}$ is considered because the solid temperature $\Theta_{1}$ is equal to the melting point, which does not change with time. Note that, as the problem is defined in the radial direction, the average temperature over the cross section could also be used instead of (10) for more accurate results. In any case, the implementation and algorithms presented later in this article are still valid for both definitions.
3 Optimal Control
3.1 Optimal control formulation
The optimal control problem for cell thawing is
[…]
The control (aka decision variable, manipulated variable, input) of this optimal control problem is the heater temperature $\Theta_{b}(\tau)$. The mechanistic model (4)–(9) describes the physics of cell thawing, and thus serves as constraints. The lower and upper bounds of the heater temperature are represented by (12).
3.2 Optimization-based approach
A detailed discussion on numerical algorithms to solve optimal control problems can be found in Nolasco et al. (2021). The common technique for solving optimal control problems is to discretize the partial differential equation (PDE) and ordinary differential equation (ODE) constraints, parameterize the time-varying control vector, and solve the resulting optimization problem numerically with a proper optimizer. This approach relies on the convergence of the optimization algorithm/solver to obtain the optimal control trajectory, and hence we denote this approach as the optimization-based approach. In this article, six different optimization-based implementations that have been used widely in optimization and optimal control applications, including on a variety of software tools and programming languages, are considered:
  • opt_Ipopt: The control vector is parameterized, and the resulting optimization is numerically solved by IPOPT (Wächter and Biegler, 2006), an open-source optimization solver for large-scale nonlinear optimization that has been implemented in various nonlinear and optimal control software packages. The IPOPT solver is employed in MATLAB via the OPTI Toolbox (Currie and Wilson, 2012). The system of ODEs is integrated by ode15s.
  • opt_fmincon: The control vector is parameterized, and the resulting optimization is numerically solved by fmincon, a built-in optimization solver for constrained nonlinear optimization in MATLAB. The system of ODEs is integrated by ode15s.
  • opt_pfmincon: The implementation is identical to opt_fmincon except that the parallel computing option for fmincon is turned on.
  • opt_sCasADi: The optimal control problem is solved using CasADi (Andersson et al., 2019), a well-established open-source tool that has been widely used for nonlinear optimization, optimal control, and model predictive control. The problem is solved symbolically by the direct single shooting method via CasADi v3.6.3, with the code directly modified from the given example pack.
  • opt_mCasADi: The implementation is similar to opt_sCasADi but uses the direct multiple shooting method instead.
  • opt_Gekko: The optimal control problem is solved using GEKKO (Beal et al., 2018), a Python package for optimization and machine learning.
3.3 Simulation-based approach
The approach in this article is inspired by a recent article on the optimal control of lithium-ion batteries (Berliner et al., 2022). The approach was motivated by the observation that the optimal control trajectories for that application moved from one active inequality constraint to another over time. That information was used to transform the optimal control problem into a mixed continuous-discrete system of index-1 DAEs in which existing software could be used to transition between active inequality constraints. This simulation-based technique eliminated the need for an optimization solver, resulting in a much more computationally efficient solution. No control vector parameterization is needed.
3.3.1 Solution of high-index DAEs
In this work, we generalize the simulation-based technique for high-index DAE systems and apply the approach to solve multiple cell thawing problems. Various numerical algorithms for solving index-1 DAEs, including the implementation used in Berliner et al. (2022), fail when the differential index is higher than 1 (Petzold, 1982), and an increase in the differential index of DAEs leads to more difficulties in obtaining the numerical solution to DAEs (Campbell et al., 2008). Many current DAE solvers are only capable of solving index-1 DAEs, so it is usually recommended to perform index reduction (e.g., reduceDAEIndex in MATLAB and dae_index_lowering in Julia) to transform a high-index system into an equivalent index-1 system, and then solve the resulting index-1 DAEs (Shampine et al., 1999; Campbell et al., 2008). This technique is usually suggested because index-1 DAEs can be solved easily by most DAE solvers, and so does not require specialized high-index DAE solvers. Nevertheless, there are two major drawbacks associated with index reduction. First, the index reduction process is costly when the number of states and differential index are huge. Second, index reduction could introduce a large number of new variables (sometimes called dummy derivatives (Mattsson and Söderlind, 1993)) to replace high-order derivatives, which unnecessarily increases the complexity of the problem. These issues are not commonly discussed in the literature as they become significant when the differential index is sufficiently large, e.g., Problem 3 in our case studies. Thus, this approach does not match our objective of accelerating the optimal control algorithm.
An alternative to the index reduction approach is to fully discretize a system of high-index DAEs (e.g., using a collocation method) and solve the discretized equations (Campbell et al., 2008). The specialized DAE solver in GEKKO employs the method of orthogonal collocation on finite elements to solve high-index DAEs (Beal et al., 2018). The main advantage of this implementation is that it does not have any constraint or limit on the differential index because index reduction and differentiation are not required (e.g., see detailed discussions by Hedengren et al. (2014) and Beal et al. (2018)). Hence, our DAE solutions for high-index systems rely on the DAE solver implemented in GEKKO.
3.3.2 Proposed algorithm
We denote this simulation-based approach as sim_DAE, which consists of three main steps. First, replace the objective function with algebraic equations, reformulating the original optimal control problem as a system of DAEs
$g(\Theta_{1}(\tau),\Theta_{2}(\tau),S(\tau),\Theta_{b}(\tau),\tau)=0.$ (13)
The choice of algebraic equations is dependent on the objective function, which is demonstrated via the case studies presented in Section 4.
Second, treat the control variable ($\Theta_{b}$ in this case) as a state instead of a decision variable. A consistent initial condition is required for this new state, and this initialization can be done in several ways, depending on the solver. For example, in GEKKO, the DAE solver does not strictly require a consistent initial condition, and so we can first guess some random initial condition and then run the DAE solver once to find the correct initial condition. Another possibility is to solve the optimization locally at the beginning of the process. Either approach requires minimal effort and computation, and so does not impact the overall computational performance. The initial condition for this new state is denoted by
$\Theta_{b}(\tau_{0})=\Theta_{b0}.$ (14)
After obtaining the DAEs and initial conditions, the final step is to solve the resulting DAEs (13) and (14). As the control is now treated as a state, the optimal control vector can be obtained directly from the DAE solver.
This DAE reformulation technique inherently assumes that at least one of the constraints, bounds, or algebraic equations resulting from reformulating the objective function is active. For many practical problems in which this strong condition does not hold, the simulation-based approach is not guaranteed to produce an optimal solution, which is discussed in more detail in Section 4.3. Changes in the active constraints or bounds can be incorporated in the simulation using a simulation-switching technique, with Section 4.4 providing a discussion that includes references to several examples.
3.4 Implementation
Both optimization- and simulation-based techniques were implemented in MATLAB R2023a following the procedures introduced in Sections 3.2 and 3.3, with GEKKO called and executed in Python 3.10. The simulations were performed on a computer equipped with an Intel® Core™ i9-13950HX Processor CPU (24 cores) and 128 GB RAM on 64-bit Windows 11. The choice and justification of important solver options are given in Appendix B.
4 Case Studies
This section presents several optimal control problems for cell thawing to demonstrate the simulation-based technique and compare it with the optimization-based techniques. These examples are drawn from real problems and protocols associated with control and optimization of cell thawing and moving boundary problems. The case studies include both simple and complex problems to help illustrate the approach and assess the relationship between the comparative performance of the various algorithms and problem complexity.
Figure 2: Trajectories for the optimal (A) heater temperature $\Theta_{b}$, (B) interface position $S$, (C) interface velocity $dS/d\tau$, (D) average temperature $\Theta_{\textrm{avg}}$, and (E) rate of temperature change $d\Theta_{\textrm{avg}}/d\tau$ in Problem 1.
4.1 Problem 1: Minimization of the thawing time
4.1.1 Problem description and formulation
This first problem is simple and intuitive, with the solution known a priori, to demonstrate the simulation-based technique and validate all of our algorithms.
In general, it has been suggested that the heating process should be done rapidly to avoid potential damages to the cells and maintain high viability (Terry et al., 2010; Baboo et al., 2019; Hunt, 2019; Uhrig et al., 2022), which is equivalent to the optimal control problem
$\min_{\Theta_{b}(\tau)}\quad\tau_{f}$ (15)
[…]
The solution to this optimal control problem is intuitive. To minimize the thawing time, the heater temperature should be fixed at its upper bound to provide the maximum heating power throughout the process.
4.1.2 DAE Reformulation
To employ the simulation-based approach, (15) needs to be reformulated as a DAE system. In this problem, the control vector $\Theta_{b}$ is treated as an algebraic state and explicitly specified an additional algebraic constraint $\Theta_{b}(\tau)=1$. As a result, the equivalent system of DAEs for (15) is
$\Theta_{b}(\tau)=1$ (16)
[…]
The number of derivatives required to transform (16) into an equivalent ODE system is 1, hence an index-1 DAE system. This index-1 DAE system can be solved easily by most DAE solvers.
4.1.3 Solution comparison
Every solution method gives the same optimal solution (Fig.  2); the heater temperature $\Theta_{b}$ is at its upper bound of 1 all the time, ensuring that the algorithms and implementation are correct. The wall-clock time required for the simulation-based technique (sim_DAE) is more than an order of magnitude lower than every optimization-based approach (Table 1). Optimization with GEKKO is the slowest method, followed by IPOPT, CasADi and fmincon. GEKKO’s ODE solvers cannot perform adaptive time-stepping, which could lead to significantly slower computation. Parallel computing fails to accelerate fmincon. The simulation-based approach is the most computationally efficient algorithm in this case study.
Table 1: Computational performance of each solution method for Problem 1.
Method Wall time (s)
opt_Ipopt $10.74\pm 0.22$
opt_fmincon $3.17\pm 0.03$
opt_pfmincon $3.23\pm 0.06$
opt_sCasADi $5.70\pm 0.02$
opt_mCasADi $5.81\pm 0.06$
opt_Gekko $34.08\pm 0.31$
sim_DAE $0.11\pm 0.01$
4.2 Problem 2: Control of the average temperature
4.2.1 Problem description and formulation
Besides using the fastest thawing as in Problem 1, some past studies considered cases in which the freezing and thawing rates are controlled, i.e., the rate of temperature change is kept constant (Seki and Mazur, 2008; Jang et al., 2017; Baboo et al., 2019; Bojic et al., 2021). As such, this problem focuses on controlling the rate of change in the average temperature during thawing by manipulating the heater temperature. The definition of the average temperature is as given in (10).
For a fixed heater temperature, the rate of temperature change is not constant (Figs.  2DE). Thus, the heater temperature needs to be manipulated. Consider the optimal control problem
[…]
where $d\Theta_{\textrm{avg}}/d\tau$ is the rate of change in the average temperature and $(d\Theta_{\textrm{avg}}/d\tau)_{\mathrm{sp}}$ is the target value (setpoint) set to 0.04 for demonstration in this case study. For the optimization-based approaches, the control vector is parameterized using a piecewise linear function, with the number of control intervals $n_{c}=16$ (Appendix A).
4.2.2 DAE Reformulation
For the simulation-based technique, first analyze the objective function in (17). This objective function is minimized when the rate of change in average temperature is equal to the setpoint. Replacing the objective function in (17) with the algebraic equation $d\Theta_{\textrm{avg}}/d\tau=(d\Theta_{\textrm{avg}}/d\tau)_{\textrm{sp}}$ results in the system of DAEs
$\dfrac{d}{d\tau}\!\!\left(\sum_{j=1}^{n}\dfrac{1}{n}{(\Theta_{2})_{j}}\right)\!=\left(\dfrac{d\Theta_{\textrm{avg}}}{d\tau}\right)_{\!\!\mathrm{sp}},$ (18)
[…]
where $\Theta_{b}$ is now treated as a state. The differential index of (18) is 2, which can be solved using GEKKO as explained in Section 3.3.
4.2.3 Solution comparison
To evaluate the correctness of the solution, define the error measured by a modified 2-norm for this problem as
$||e||_{2}=\sqrt{\dfrac{1}{n_{k}}\sum_{k=1}^{n_{k}}\left(\left(\dfrac{d\Theta_{\textrm{avg}}}{d\tau}\right)_{\!k}-\left(\dfrac{d\Theta_{\textrm{avg}}}{d\tau}\right)_{\!\!\mathrm{sp}}\right)^{\!\!2}},$ (19)
where $(d\Theta_{\textrm{avg}}/d\tau)_{k}$ is the rate of change in average temperature resulting from solving (17) or (18) evaluated at each time point $k$ in the time span [$\tau_{0},\tau_{f}$] and $n_{k}$ is the number of sampling time points. Small $||e||_{2}$ corresponds to the rate of change in average temperature being close to the target value, indicating that the heater temperature is optimal; i.e., the solution method is accurate.
The optimal solutions to Problem 2 obtained from the optimization- and simulation-based approaches are illustrated in Fig.  3. Most solution techniques predict the same heater temperature profile except the methods with fmincon that predict a slightly higher temperature (Fig.  3A). As time progresses, the heater temperature increases to compensate for a reduction in the temperature difference (driving force) between the heater and product. The rate of change in average temperature can be controlled nearly perfectly at 0.04 for IPOPT, CasADi, GEKKO, and the simulation-based method, while this value fluctuates up to about 0.046 for fmincon, implying that fmincon fails to converge to the optimal solution in this case (Fig.  3B). It is also evident that the average temperature increases linearly at the rate of 0.04 in all cases except for fmincon (Fig.  3C).
Figure 3: Trajectories for the optimal (A) heater temperature, (B) rate of change in average temperature, and (C) average temperature simulated via the optimization- and simulation-based approaches in Problem 2.
Table 2: Computational performance and accuracy of each solution method for Problem 2.
Method Wall time (s) $||e||_{2}$
opt_Ipopt $162.67\pm 3.72$ 8.03×$\times 10^{-5}$
opt_fmincon $349.07\pm 11.67$ 1.21×$\times 10^{-3}$
opt_pfmincon $79.40\pm 1.57$ 1.21×$\times 10^{-3}$
opt_sCasADi $363.36\pm 12.52$ 8.04×$\times 10^{-5}$
opt_mCasADi $5189.15\pm 71.42$ 8.04×$\times 10^{-5}$
opt_Gekko $191.89\pm 4.86$ 4.91×$\times 10^{-5}$
sim_DAE $3.71\pm 0.15$ 2.25×$\times 10^{-5}$
The simulation-based approach is by far the fastest solution method, accelerating the computation by 40$\times$ to 90$\times$ compared to the typical solvers IPOPT and fmincon (Table 2). Without parallel computing, optimization with IPOPT is fastest, followed by GEKKO, fmincon, and CasADi. Parallel computing can significantly reduce the computation time for fmincon. With the direct multiple shooting method (opt_mCasADi), the number of states resulting from spatial discretization is large. For example, with $n=20$, there are 41 states for each control interval and 656 states for $n_{c}=16$. This large nonlinear optimization is computationally expensive to solve compared to other approaches. In terms of accuracy, fmincon produces the largest error, which is about 30–40 fold higher than for the other techniques. This indicates that the solution obtained from fmincon is not optimal, agreeing with the plots shown in Fig.  3. Although both fmincon and IPOPT employ the interior point methods, the detailed implementations (e.g., gradient/hessian calculation, scaling) are not exactly the same, which could result in different accuracy and performance (see detailed discussion and comparison of these solvers in Rojas-Labanda and Stolpe (2015)). The simulation-based and other optimization-based methods except fmincon have similar accuracy. Despite an increase in the differential index, the simulation-based approach is still much faster than the optimization-based techniques while giving the same level of accuracy.
4.3 Problem 3: Control of the interface velocity
4.3.1 Problem description and formulation
One of the primary interests in a Stefan problem is the moving solid-liquid interface. The literature has studied and shown the importance of interface position tracking and optimization in various applications (Hill and Wu, 1994; Brezina et al., 2018; Srisuma et al., 2023b, for example). This problem focuses on the control of the interface velocity, i.e., the evolution of a melting/freezing process.
Consider the optimal control problem
[…]
where $dS/d\tau$ is the interface velocity and $\left(dS/d\tau\right)_{\mathrm{sp}}$ is the target velocity set to $-0.1$ for demonstration in this case study. In Problem 1 (Fig.  2BC), the velocity of the moving interface is not constant if the heater temperature is fixed. In this case study, the aim is to control the velocity of the moving interface to be constant at $(dS/d\tau)_{\mathrm{sp}}$ by manipulating the heater temperature throughout the process as formulated in (20). For the optimization-based approaches, the control vector is parameterized using a piecewise linear function, with the number of control intervals $n_{c}=12$ (see Appendix A).
4.3.2 DAE Reformulation
To apply the simulation-based technique, a similar approach used in Problem 2 can be used, but here we instead enforce the moving interface velocity by modifying the interface equation (6) to be equal to $(dS/d\tau)_{\textrm{sp}}$. Consequently, (20) is reformulated as the DAE system
$\dfrac{dS}{d\tau}=f_{3}(\Theta_{1},\Theta_{2},S)=\left(\dfrac{dS}{d\tau}\right)_{\!\!\mathrm{sp}},$ (21)
[…]
The differential index of (21) is $n$, which is the number of grid points resulting from spatial discretization. Therefore, if the spatial discretization of the domain is made finer to increase numerical accuracy, the differential index of (21) increases. In this case, $n=20$, and hence (21) becomes an index-20 DAE system, making the problem even more complicated and difficult to solve compared to the index-2 system in Problem 2.
To explore the possibility of using the index reduction technique, MATLAB’s reduceDAEIndex was used to reduce the differential index of (21). However, the index reduction process was not complete even after 20 hours of simulation. Obviously, this technique is not feasible for our problem as mentioned before in Section 3.3. Therefore, GEKKO’s DAE solver is needed.
4.3.3 Solution comparison
Similar to Problem 2, we define the error measured by a modified 2-norm for this problem as
$||e||_{2}=\sqrt{\dfrac{1}{n_{k}}\sum_{k=1}^{n_{k}}\left(\left(\dfrac{dS}{d\tau}\right)_{\!\!k}-\left(\dfrac{dS}{d\tau}\right)_{\!\!\mathrm{sp}}\right)^{\!\!2}},$ (22)
where $(dS/d\tau)_{k}$ is the interface velocity resulting from solving (20) or (21) evaluated at each time point $k$ in the time span [$\tau_{0},\tau_{f}$].
The optimal solutions to Problem 3 are shown in Fig.  4. The optimal heater temperature predicted by each technique is nearly identical except that there is a large drop around the end of the process for the fmincon cases (Fig.  4A). The interface velocity can be controlled at about $-0.1$ all the time for IPOPT, CasADi, GEKKO, and the simulation-based technique, while some fluctuation is observed for fmincon (Fig.  4B). This indicates that fmincon is less accurate than the others approaches, which is a similar trend observed before in Problem 2. With the interface velocity controlled, the interface position recedes linearly at the rate of 0.1 (Fig.  4C).
Figure 4: Trajectories for the optimal (A) heater temperature, (B) interface velocity, and (C) interface position via the optimization- and simulation-based approaches in Problem 3.
For the computational performance, the simulation-based technique is about 60$\times$ to 900$\times$ faster than all of the optimization-based approaches (Table 3). IPOPT, GEKKO, and CasADi are comparable for the direct single shooting method, while the multiple shooting method is much slower. Parallel computing can significantly accelerate fmincon but is still much slower than the simulation-based approach. fmincon is the least accurate method; the error is higher than for the other approaches by more than an order of magnitude. Except for fmincon, the simulation-based approach has similar accuracy as the optimization-based approaches, with the error of about $2$$\times$$10^{-5}$.
Table 3: Computational performance and accuracy of each solution method for Problem 3.
Method Wall time (s) $||e||_{2}$
opt_Ipopt $214.01\pm 5.75$ 1.96×$\times 10^{-5}$
opt_fmincon $317.89\pm 5.46$ 3.05×$\times 10^{-4}$
opt_pfmincon $92.30\pm 1.52$ 3.05×$\times 10^{-4}$
opt_sCasADi $225.42\pm 5.92$ 1.96×$\times 10^{-5}$
opt_mCasADi $2618.13\pm 83.29$ 1.97×$\times 10^{-5}$
opt_Gekko $179.77\pm 2.45$ 4.37×$\times 10^{-5}$
sim_DAE $2.99\pm 0.07$ 2.41×$\times 10^{-5}$
From the three case studies, the simulation-based approach reliably solves the optimal control problems irrespective of the differential index, ranging from a simple index-1 problem to an extremely high-index DAE system. The approach is much faster than any optimization-based technique while giving the same level of accuracy.
The simulation-based approach, however, has one limitation associated with the DAE reformulation. As described in Section 3.3, the DAE reformulation is guaranteed to be optimal when at least one of the constraints, bounds, or algebraic equations resulting from reformulating the objective function is active. This is typically the case when the objective function of the original optimal control problem can be minimized by solving each subproblem resulting from control vector parameterization individually – in other words, local minimization at every control/time interval (subproblem) is equivalent to global minimization of the whole process (original problem). For example, in Problem 1, if the thawing time of every control interval is minimized, the total thawing time is automatically minimized. In Problems 2 and 3, if the rate of temperature change and interface velocity are controlled locally for each control interval, the whole process is also controlled. These objective functions are common in many applications, for which the simulation-based technique can be applied, including lithium-ion batteries of various chemistries (Berliner et al., 2022; Galuppini et al., 2023; Matschek et al., 2023) and microwave lyophilization (Srisuma et al., 2023a). An example of a system in which the method is not optimal is a reactor with multiple chemical reactions, with the goal of maximizing the amount of desired product at the end of the process. Maximizing the amount of product during the first control interval could concurrently increase the amount of some byproducts that can degrade the desired product later. In this situation, local optimization is not equivalent to global optimization. As such, an algebraic equation/constraint may not be active, so an optimization-based technique is needed. Although the simulation-based approach does not give an optimal solution to such problems, it can be used to provide an initialization to the optimization solver.
4.4 Problem 4: Sensitivity analysis
In Problems 2 and 3, the simulation-based approach is tested with one value of $(d\Theta_{\textrm{avg}}/d\tau)_{\mathrm{sp}}$ and $(dS/d\tau)_{\textrm{sp}}$, which are 0.04 and $-0.1$, respectively. To demonstrate the robustness of our framework, this section conducts a sensitivity analysis by varying $(d\Theta_{\textrm{avg}}/d\tau)_{\mathrm{sp}}$ and $(dS/d\tau)_{\textrm{sp}}$ and employs the simulation-based approach to solve the problems.
The simulation-based approach provides accurate solutions irrespective of the values of $(d\Theta_{\textrm{avg}}/d\tau)_{\mathrm{sp}}$ and $(dS/d\tau)_{\textrm{sp}}$ (Figs.  5 and 6). The rate of change in average temperature and interface velocity are at the setpoints except for $(dS/d\tau)_{\textrm{sp}}=-0.15$. An interface velocity of $-0.15$ is too fast for the given heater temperature, so the upper bound is active, and that the heater temperature cannot be increased further to achieve the target velocity. The target velocity that is larger (smaller in magnitude) than the peak of Fig.  2C, which about $-0.12$, will never violate the bound (see Appendix C for a formal proof).
For cases where the bound is active, a simulation switching technique can be used to transition between active constraints, resulting in a hybrid discrete/continuous dynamic simulation (Feehery and Barton, 1998; Barton et al., 2000; Berliner et al., 2022). First, initialize a DAE simulation as usual. Next, terminate the current simulation when the bound on any control vector is active and then initialize a new DAE simulation with that active bound/constraint. Perform this switching whenever there is a change in the active constraints, which results in a mixed continuous-discrete DAE system instead of a pure continuous system. The complexity of this procedure depends on the choice of a DAE solver. For example, MATLAB’s and Julia’s DAE solvers have a built-in function to track all variables and terminate a simulation when the specified condition (aka event) is met, which facilitates the implementation of such switching. The switching technique can also be used to handle cases where the control trajectory is discontinuous (see examples in Berliner et al. (2022); Srisuma et al. (2023a)). GEKKO’s DAE solvers, on the other hand, do not have a built-in option for handling switches, and so this process needs to be executed manually (see our software/code for implementation). This switching technique can also be used for path constraints.
Figure 5: Trajectories for the optimal (A) heater temperature, (B) rate of change in average temperature, and (C) average temperature simulated via the simulation-based approach at four different values of $(d\Theta_{\textrm{avg}}/d\tau)_{\mathrm{sp}}$ in Problem 4.
Figure 6: Trajectories for the optimal (A) heater temperature, (B) interface velocity, and (C) interface position simulated via the simulation-based approach at four different values of $(dS/d\tau)_{\textrm{sp}}$ in Problem 4.
With this capability, the benefit of the simulation-based approach is even more substantial in applications where a large number of simulation runs are required, e.g., parametric studies and design optimization.
5 Conclusion
This article describes a new approach for determining control and state trajectories in optimal control problems by reformulation as a system of DAEs. The optimal control and state vectors are obtained via simulation of the resulting DAE system with the selected DAE solver, eliminating the need for an optimization solver and thus greatly accelerating the computation.
Our proposed framework was demonstrated and benchmarked against a variety of optimization-based approaches – IPOPT, fmincon, CasADi, and GEKKO – for case studies involving the optimal control of a Stefan problem for cell thawing. The case studies include the minimization of the thawing time, the control of the average temperature, and the control of the interface velocity, transformed into DAE systems of index 1, 2, and 20, respectively. At the same level of accuracy, the simulation-based approach is more than an order of magnitude faster than every optimization-based method in all cases. The simulation-based approach offers an efficient and reliable framework for solving the optimal control problems, serving as a promising alternative to the traditional techniques in applications where speed is crucial, e.g., real-time online model predictive control.
Some future directions of interest are to (1) generalize the approach and (2) improve the computational performance. The ultimate goal is a highly efficient optimal control solver applicable to problems with a variety of objective functions and constraints that arise in multiple fields. Implementation would be simplified if high-index DAE solvers become available that have built-in options for mixed continuous-discrete simulation.
Data Availability
All software and data will be made available upon publication of the manuscript.
Acknowledgments
The authors would like to thank Dr. John D. Hedengren for advice on the software GEKKO.
Appendices
Appendix A Control Vector Parameterization
A.1 Piecewise constant and linear controls
Piecewise constant and linear controls are most common in optimal control (Nolasco et al., 2021). This section justifies the use of piecewise linear controls in Problems 2 and 3 for the optimization-based approaches, in comparison to piecewise constant controls. This comparison considers the most complex problem (Problem 3), with opt_Ipopt as the solver.
Table 4: Comparison between the errors $||e||_{2}$ resulting from piecewise constant and linear control vector parameterization for different values of control intervals $n_{c}$.
$n_{c}$ Error measured by $||e||_{2}$
Piecewise constant Piecewise linear
4 0.012 3.01$\times$$10^{-4}$
8 0.010 5.68$\times$$10^{-5}$
12 0.012 1.96$\times$$10^{-5}$
From Table 4, the piecewise linear control is more accurate than the piecewise constant control by many orders of magnitude. Since the dynamics of the interface position and average temperature are highly nonlinear, a large value of $n_{c}$ is required for a piecewise constant control to accurately manipulate the average temperature and moving interface, leading to a much larger nonlinear program that is computationally expensive and difficult to converge. As a result, a piecewise linear control is selected in this work.
A.2 Number of control intervals
The accuracy of the optimization-based solution is influenced by the number of control intervals $n_{c}$. This section justifies those numbers by investigating the effect of $n_{c}$ on the accuracy of the solution.
The optimal control problems defined by (17) and (20) were solved via opt_Ipopt and $||e||_{2}$ was calculated for different values of $n_{c}$. Logically, the accuracy of the solution should improve, i.e., smaller $||e||_{2}$, when $n_{c}$ increases due to finer discretization. From Fig.  7, the accuracy of the solution does not change significantly after $n_{c}$ reaches some certain value, and so this threshold should be chosen for control vector parameterization. As a result, $n_{c}$ was set to 16 for Problem 2 (Fig.  7A) and 12 for Problem 3 (Fig.  7B).
Figure 7: Errors measured by $||e||_{2}$ for different values of control intervals $n_{c}$ for (A) Problem 2 and (B) Problem 3.
Appendix B Solver Options
There are a large number of options for optimization, ODE, and DAE solvers, which could lead to different accuracy and computational performance. This section describes and justifies the choice of important solver options used in this work. The default values were used for the options not mentioned here.
For all case studies, index-1 DAEs were solved by ode15s in MATLAB, whereas high-index DAEs were handled by GEKKO in Python. The differential index of DAEs was checked using reduceDAEIndex. The wall-clock times (aka wall times, clock times) were measured using the tic and toc functions, with each simulation repeated at least 10 times for consistent results, as indicated by the standard deviation of the measured wall times being smaller than 5%. For fair comparison of computational performance, the initial guess of $\Theta_{b}$ was set to the average value of 0.5 in all cases.
For the optimization-based approaches, the optimality tolerance was set to $10^{-7}$ for all optimization solvers, including IPOPT, fmincon, CasADi, and GEKKO. The integration tolerance for ode15s was set to $10^{-10}$. Using higher values for the optimality and integration tolerances could lead to inaccurate solutions, while using tighter values does not improve the accuracy significantly. For the objective functions (17) and (20) in Problems 2 and 3, the derivatives were approximated using a finite difference scheme, whereas the integrals can be approximated using a Riemann sum. The time step $\Delta\tau$ was set to 0.05.
While ode15s uses adaptive time-stepping, this option is not available in GEKKO; i.e., users have to specify a vector of time points for ODE integration. Hence, there is no ideal comparison between GEKKO’s ODE solvers and MATLAB’s ode15s. The fixed time step $\Delta\tau=0.05$ was specified for ODE integration in GEKKO, with five collocation points (INODES = 5) for each time interval. This time step was chosen to be consistent with the value used for a finite difference and Riemann sum approximation mentioned in the previous paragraph, while the number of collocation points was selected such that that the temperature and interface position simulated by GEKKO’s ODE solver have the same level of accuracy as those simulated by ode15s.
For the simulation-based approach, ode15s was used for an index-1 system (Problem 1), with the exact same solver options as mentioned above. GEKKO’s DAE solver was used for high-index systems (Problems 2 and 3). As a reformulated DAE system is different from the original ODE system, the time step needs to be adjusted accordingly. The time step was selected such that the accuracy of the simulation-based solution measured by $||e||_{2}$ is on the same order of magnitude as that of the optimization-based solution, giving $\Delta\tau=0.12$ and $\Delta\tau=0.37$ for Problems 2 and 3, respectively.
The accuracy of both the optimization- and simulation-based approaches is measured using $||e||_{2}$ defined by (19) and (22). To calculate $||e||_{2}$, the derivatives were approximated using a finite difference scheme with $\Delta\tau=0.05$.
Appendix C Bounds on the Interface Velocity and Rate of Change in Temperature
We show in Fig.  2, Problems 2 and 3, and Section 4.4 that, if the target velocity (and also the rate of temperature change) is chosen properly, the bound on a control vector will never be violated. A formal proof is given in this section.
To obtain a closed-form solution for the proof, the heat conduction in the liquid domain is assumed to be quasi-steady. This approximation is accurate for problems associated with phase change as most of the heat transfer occurs at the moving interface. In this case, the closed-form solution for the interface velocity is
$\frac{dS}{d\tau}=-\frac{k_{2}U\Theta_{b}(T_{0}-T_{m})}{\rho L\alpha_{1}S\left(U\ln(1/S)+k/b\right)},$ (23)
where all the parameter description and values can be found in Srisuma et al. (2023b). From (23), the interface velocity $(dS/d\tau)$ is a monotonic function of the heater temperature $(\Theta_{b})$. Hence, the condition in which the bounds on the heater temperature are not violated can be obtained by substituting the upper bound $\Theta_{b}=1$ into (23), which is
$\left(\dfrac{dS}{d\tau}\right)_{\!\!\mathrm{sp}}>-\frac{k_{2}U(T_{0}-T_{m})}{\rho\alpha_{1}\Delta H_{\!f}S\left(U\ln(1/S)+k/b\right)}.$ (24)
With the parameter values in Srisuma et al. (2023b), the right-hand side of (24) is about $-0.13$. This conclusion agrees with the results presented in Problems 3 and 4, in which the bounds are not violated when $(dS/d\tau)_{\mathrm{sp}}$ is higher than $-0.13$, whereas the upper bound is active when $(dS/d\tau)_{\mathrm{sp}}$ is $-0.15$. A similar strategy can be used for analyzing the temperature change, e.g., using a thermal lumped capacity approximation.
Note that our simulation-based approach does not require to know a priori if the bounds will be active or not, as the upper and lower bounds can be handled as shown in Problem 4.

### v2

@@
 Abstract
-This article describes a new, efficient way of finding control and state trajectories in optimal control problems by transformation into a system of differential-algebraic equations (DAEs).
+This article describes a new, efficient way of finding control and state trajectories in optimal control problems by reformulation as a system of differential-algebraic equations (DAEs).
 The optimal control and state vectors can be obtained via simulation of the resulting DAE system with the selected DAE solver, eliminating the need for an optimization solver.
-Our simulation-based framework was demonstrated and benchmarked against various optimization-based approaches via four case studies associated with optimization and control of a Stefan problem with application to cell therapy.
-The simulation-based approach is faster than every optimization-based method by more than an order of magnitude while giving the same level of accuracy in all cases.
+Our simulation-based approach is demonstrated and benchmarked against various optimization-based algorithms via four case studies associated with the optimization and control of a Stefan problem for cell therapy.
+The simulation-based approach is faster than every optimization-based method by more than an order of magnitude while giving similar/better accuracy in all cases.
+The solution obtained from the simulation-based approach is guaranteed to be optimal provided that at least one constraint or algebraic equation resulting from the reformulation remains active at all times.
 The proposed technique offers an efficient and reliable framework for optimal control, serving as a promising alternative to the traditional techniques in applications where speed is crucial, e.g., real-time online model predictive control.
@@
 Numerical techniques have been developed for implementing and simulating the Stefan problems (Meyer, 1971; Velardi and Barresi, 2008; Kurbatova and Ermolaeva, 2019; Gusev et al., 2021; Srisuma et al., 2023b).
-Optimal control of Stefan problems was extensively investigated and proven useful for various industrial applications over the past few decades, with many different objective functions, constraints, and controls (manipulated variables) considered.
-Some examples are the control of the heating process to satisfy the heating speed and thermoelastic stress constraints by manipulating the furnace temperature (Roubíček, 1986), the stabilization of the moving boundary and temperature/concentration fields by varying the heat flux (Pawlow, 1987), the maximization of the amount of melted solid in the melting process via controlling the heat flux (Silva Neto and White, 1994), the control of the moving boundary to follow the desired path in the solidification process by manipulating the wall temperature (Hinze and Ziegenbalg, 2007), the control of the water level in the drainage basin by varying the discharge velocity (Miyaoka and Kawahara, 2008), the optimization of a thin-film drying process via manipulating the air temperature (Mesbah et al., 2014), and the minimization of the metallurgical length (ML) deviation in steel casting via controlling the boundary heat flux (Chen et al., 2019).
+Optimal control of Stefan problems has been extensively investigated and proven useful for various industrial applications over the past few decades, with many different objective functions, constraints, and controls (manipulated variables) considered.
+Some examples are the control of a heating process to satisfy constraints on the heating speed and thermoelastic stress by manipulating the furnace temperature (Roubíček, 1986), the stabilization of the moving boundary and temperature/concentration fields by varying the heat flux (Pawlow, 1987), the maximization of the amount of melted solid in a melting process via controlling the heat flux (Silva Neto and White, 1994), the control of the moving boundary to follow the desired path in a solidification process by manipulating the wall temperature (Hinze and Ziegenbalg, 2007), the control of the water level in a drainage basin by varying the discharge velocity (Miyaoka and Kawahara, 2008), the optimization of a thin-film drying process via manipulating the air temperature (Mesbah et al., 2014), and the minimization of the metallurgical length deviation in steel casting by controlling the boundary heat flux (Chen et al., 2019).
 All the aforementioned studies rely on optimization algorithms/solvers to obtain the optimal control trajectories.
@@
 The PDAE system is then solved numerically by feeding the differential-algebraic equations (DAEs) obtained by spatial discretization into a DAE solver.
-In this approach, the optimal control vector is obtained directly by a DAE solver, without using any optimization solver, resulting in a highly computationally efficient solution to the optimal control problem.
-Recently, this approach was used to solve optimal control problems associated with a Stefan problem (Srisuma et al., 2024).
-Although optimal control with Stefan problems has been explored for a wide range of processes, applications to cell therapy have not been available.
-Previous studies showed that accurate prediction, control, and optimization of cryopreservation and cell thawing can improve the viability and quality of the resulting cells, which directly benefits cell therapy (Seki and Mazur, 2008; Jang et al., 2017; Baboo et al., 2019; Hunt, 2019; Bojic et al., 2021; Cottle et al., 2022; Uhrig et al., 2022).
+In this approach, the optimal control vector is obtained directly from the DAE solver, without using any optimization solver, resulting in a highly computationally efficient solution to the optimal control problem.
+Recently, this approach was used in the optimal control of a Stefan problem (Srisuma et al., 2024).
+Although optimal control of Stefan problems has been explored for a wide range of processes, applications to cell therapy are lacking.
+Previous studies have shown that accurate prediction, control, and optimization of cryopreservation and cell thawing can improve the viability and quality of the resulting cells, which directly benefits cell therapy (Seki and Mazur, 2008; Jang et al., 2017; Baboo et al., 2019; Hunt, 2019; Bojic et al., 2021; Cottle et al., 2022; Uhrig et al., 2022).
 These benefits and case studies therefore motivate the development of an efficient Stefan problem-based optimal control algorithm for cell therapy.
@@
   1. 1.
-    derive the simulation-based approach for optimal control of a Stefan problem for the cell thawing process in cell therapy;
+    derive the simulation-based approach for optimal control of a Stefan problem applied to the cell thawing process in cell therapy;
   2. 2.
-    generalize the simulation-based approach for both index-1 and high-index DAE systems;
+    generalize the simulation-based approach to both index-1 and high-index DAE systems;
   3. 3.
-    demonstrate and benchmark the simulation-based algorithm against various optimization-based optimal control algorithms; and
+    demonstrate and benchmark the simulation-based algorithm against various optimization-based optimal control algorithms;
   4. 4.
-    apply the proposed approach to efficiently and accurately solve several practical problems associated with optimization and control of cell thawing.
+    apply the proposed approach to efficiently and accurately solve several practical problems associated with the optimization and control of cell thawing; and
+  5. 5.
+    discuss its limitation and condition required for the approach to provide an optimal solution.
 This article is organized as follows.
@@
 2 Stefan Problem and Cell Thawing
-The system used for the demonstration of our optimal control technique in this article concerns cell thawing, which is a process used in cell therapy before cells are introduced to the patients (Fig.
+The system used for the demonstration of our optimal control technique in this article concerns cell thawing, which is a process used in cell therapy before cells are introduced into the patients (Fig.
 1).
@@
 All variables are written in dimensionless form.
-In dimensionless form, the model consists of three main parts which describe the evolution of the solid temperature (4), the liquid temperature (5), and the interface position (6), respectively.
-The energy is continuously supplied to the system by the heater with the temperature $\Theta_{b}$.
-The system initially consists of pure solid, which corresponds to $S=1$; i.e., the interface position is at the outer boundary.
-As the thawing progresses, the interface position decreases and eventually becomes 0, indicating complete melting or pure liquid.
-The maximum temperature (dimensionless) is equal to 1, which is equivalent to $37\,^{\circ}$, the maximum temperature recommended for cell thawing (Baboo et al., 2019; Uhrig et al., 2022).
-The minimum temperature (dimensionless) is equal to 0, corresponding to the melting point of $-2\,^{\circ}$.
+In dimensionless form, the model consists of three main parts that describe the evolution of the solid temperature (4), the liquid temperature (5), and the interface position (6), respectively.
+The system is initially in the solid state, which corresponds to $S=1$; i.e., the interface position is at the outer boundary.
+As the thawing progresses, the interface position decreases and eventually becomes 0, indicating complete melting.
+The heater temperature is $\Theta_{b}$, which can be fixed or varied with time.
+The maximum dimensionless temperature is 1, which is equivalent to $37\,^{\circ}$, the highest temperature recommended for cell thawing (Baboo et al., 2019; Uhrig et al., 2022).
+The minimum dimensionless temperature is 0, corresponding to the melting point of $-2\,^{\circ}$.
 Finally, since the temperature varies spatially, we define the average temperature $\Theta_{\textrm{avg}}$ as
@@
 In this case, only the liquid temperature $\Theta_{2}$ is considered because the solid temperature $\Theta_{1}$ is equal to the melting point, which does not change with time.
-Note that, as the problem is defined in the radial direction, the average temperature over the cross section could also be used instead of (10) for more accurate results.
+Note that, as our problem is defined in the radial direction, the average temperature over the cross section could also be used instead of (10) for more accurate results.
 In any case, the implementation and algorithms presented later in this article are still valid for both definitions.
@@
 The control (aka decision variable, manipulated variable, input) of this optimal control problem is the heater temperature $\Theta_{b}(\tau)$.
-The mechanistic model (4)–(9) describes the physics of cell thawing, and thus serves as constraints.
-The lower and upper bounds of the heater temperature are represented by (12).
+The mechanistic model (4)–(9) describes the physics of cell thawing and thus serves as the constraints of the optimization.
+The lower and upper bounds on the heater temperature are represented by (12).
 3.2 Optimization-based approach
-A detailed discussion on numerical algorithms to solve optimal control problems can be found in Nolasco et al.
+A detailed discussion on numerical algorithms for solving optimal control problems can be found in Nolasco et al.
 (2021).
-The common technique for solving optimal control problems is to discretize the partial differential equation (PDE) and ordinary differential equation (ODE) constraints, parameterize the time-varying control vector, and solve the resulting optimization problem numerically with a proper optimizer.
+The common technique for solving optimal control problems is to discretize the partial differential equation (PDE) and ordinary differential equation (ODE) constraints, parameterize the time-varying control vector, and solve the resulting optimization numerically with a proper optimizer.
 This approach relies on the convergence of the optimization algorithm/solver to obtain the optimal control trajectory, and hence we denote this approach as the optimization-based approach.
-In this article, six different optimization-based implementations that have been used widely in optimization and optimal control applications, including on a variety of software tools and programming languages, are considered:
+In this article, six different optimization-based implementations that have been used widely in optimization and optimal control applications, including a variety of software tools and programming languages, are considered:
   • opt_Ipopt: The control vector is parameterized, and the resulting optimization is numerically solved by IPOPT (Wächter and Biegler, 2006), an open-source optimization solver for large-scale nonlinear optimization that has been implemented in various nonlinear and optimal control software packages.
 The IPOPT solver is employed in MATLAB via the OPTI Toolbox (Currie and Wilson, 2012).
-The system of ODEs is integrated by ode15s.
+The system of ODEs is integrated using ode15s.
   • opt_fmincon: The control vector is parameterized, and the resulting optimization is numerically solved by fmincon, a built-in optimization solver for constrained nonlinear optimization in MATLAB.
-The system of ODEs is integrated by ode15s.
+The system of ODEs is integrated using ode15s.
   • opt_pfmincon: The implementation is identical to opt_fmincon except that the parallel computing option for fmincon is turned on.
@@
 The approach was motivated by the observation that the optimal control trajectories for that application moved from one active inequality constraint to another over time.
-That information was used to transform the optimal control problem into a mixed continuous-discrete system of index-1 DAEs in which existing software could be used to transition between active inequality constraints.
-This simulation-based technique eliminated the need for an optimization solver, resulting in a much more computationally efficient solution.
+That information was used to transform the optimal control problem into a mixed continuous-discrete system of index-1 DAEs, in which existing software can be used to transition between active inequality constraints.
+This simulation-based technique eliminates the need for an optimization solver, resulting in a much more computationally efficient solution.
 No control vector parameterization is needed.
 3.3.1 Solution of high-index DAEs
-In this work, we generalize the simulation-based technique for high-index DAE systems and apply the approach to solve multiple cell thawing problems.
+In this work, we extend the simulation-based technique to handle high-index DAE systems and apply the approach to solve multiple cell thawing problems.
 Various numerical algorithms for solving index-1 DAEs, including the implementation used in Berliner et al.
-(2022), fail when the differential index is higher than 1 (Petzold, 1982), and an increase in the differential index of DAEs leads to more difficulties in obtaining the numerical solution to DAEs (Campbell et al., 2008).
+(2022), fail when the differential index is higher than 1 because an increase in the differential index of DAEs introduces more difficulties in obtaining numerical solutions (Petzold, 1982; Campbell et al., 2008).
 Many current DAE solvers are only capable of solving index-1 DAEs, so it is usually recommended to perform index reduction (e.g., reduceDAEIndex in MATLAB and dae_index_lowering in Julia) to transform a high-index system into an equivalent index-1 system, and then solve the resulting index-1 DAEs (Shampine et al., 1999; Campbell et al., 2008).
-This technique is usually suggested because index-1 DAEs can be solved easily by most DAE solvers, and so does not require specialized high-index DAE solvers.
+This technique is generally preferable because index-1 DAEs can be solved easily by most DAE solvers, eliminating the need for specialized high-index DAE solvers.
 Nevertheless, there are two major drawbacks associated with index reduction.
 First, the index reduction process is costly when the number of states and differential index are huge.
-Second, index reduction could introduce a large number of new variables (sometimes called dummy derivatives (Mattsson and Söderlind, 1993)) to replace high-order derivatives, which unnecessarily increases the complexity of the problem.
+Second, index reduction could introduce a large number of new variables, sometimes called dummy derivatives (Mattsson and Söderlind, 1993), to replace high-order derivatives, which unnecessarily increases the complexity of the problem.
 These issues are not commonly discussed in the literature as they become significant when the differential index is sufficiently large, e.g., Problem 3 in our case studies.
@@
 The specialized DAE solver in GEKKO employs the method of orthogonal collocation on finite elements to solve high-index DAEs (Beal et al., 2018).
-The main advantage of this implementation is that it does not have any constraint or limit on the differential index because index reduction and differentiation are not required (e.g., see detailed discussions by Hedengren et al.
+The main advantage of this implementation is that it does not entail any index reduction or dummy derivatives, making it more efficient when dealing with some high-index problems (e.g., see detailed discussions by Hedengren et al.
 (2014) and Beal et al.
@@
 Second, treat the control variable ($\Theta_{b}$ in this case) as a state instead of a decision variable.
-A consistent initial condition is required for this new state, and this initialization can be done in several ways, depending on the solver.
-For example, in GEKKO, the DAE solver does not strictly require a consistent initial condition, and so we can first guess some random initial condition and then run the DAE solver once to find the correct initial condition.
+A consistent initial condition is required for this new state, and the initialization can be done in several ways, depending on the solver.
+For example, in GEKKO, the DAE solver does not strictly require a consistent initial condition, so it is possible to start with some reasonable guess and then run the DAE solver once to find a more accurate initial condition.
 Another possibility is to solve the optimization locally at the beginning of the process.
-Either approach requires minimal effort and computation, and so does not impact the overall computational performance.
+Either approach requires minimal effort and computation, and therefore does not impact the overall computational performance.
 The initial condition for this new state is denoted by
@@
 After obtaining the DAEs and initial conditions, the final step is to solve the resulting DAEs (13) and (14).
-As the control is now treated as a state, the optimal control vector can be obtained directly from the DAE solver.
-This DAE reformulation technique inherently assumes that at least one of the constraints, bounds, or algebraic equations resulting from reformulating the objective function is active.
-For many practical problems in which this strong condition does not hold, the simulation-based approach is not guaranteed to produce an optimal solution, which is discussed in more detail in Section 4.3.
-Changes in the active constraints or bounds can be incorporated in the simulation using a simulation-switching technique, with Section 4.4 providing a discussion that includes references to several examples.
+As the control is now treated as a state, the optimal control vector is obtained automatically from the DAE solver.
+This DAE reformulation technique inherently assumes that at least one of the inequality constraints, bounds, or algebraic equations resulting from reformulating the objective function is active.
+For problems in which this condition does not hold, the simulation-based approach is not guaranteed to produce an optimal solution, which is discussed in more detail in Section 4.3.
+Changes in the active constraints or bounds can be incorporated in the simulation using a switching technique, with Section 4.4 providing a discussion that includes references to several examples.
 3.4 Implementation
-Both optimization- and simulation-based techniques were implemented in MATLAB R2023a following the procedures introduced in Sections 3.2 and 3.3, with GEKKO called and executed in Python 3.10.
-The simulations were performed on a computer equipped with an Intel® Core™ i9-13950HX Processor CPU (24 cores) and 128 GB RAM on 64-bit Windows 11.
-The choice and justification of important solver options are given in Appendix B.
+In this work, both optimization- and simulation-based techniques were implemented in MATLAB R2023a following the procedures introduced in Sections 3.2 and 3.3, with GEKKO called and executed in Python 3.10.
+All simulations were performed on a computer equipped with an Intel® Core™ i9-13950HX processor (24 cores) and 128 GB RAM running 64-bit Windows 11.
+The choice and justification of important solver options are given in Appendix A.
 4 Case Studies
@@
 The case studies include both simple and complex problems to help illustrate the approach and assess the relationship between the comparative performance of the various algorithms and problem complexity.
-Figure 2: Trajectories for the optimal (A) heater temperature $\Theta_{b}$, (B) interface position $S$, (C) interface velocity $dS/d\tau$, (D) average temperature $\Theta_{\textrm{avg}}$, and (E) rate of temperature change $d\Theta_{\textrm{avg}}/d\tau$ in Problem 1.
+Figure 2: Trajectories of the optimal (A) heater temperature $\Theta_{b}$, (B) interface position $S$, (C) interface velocity $dS/d\tau$, (D) average temperature $\Theta_{\textrm{avg}}$, and (E) rate of temperature change $d\Theta_{\textrm{avg}}/d\tau$ for Problem 1.
 4.1 Problem 1: Minimization of the thawing time
 4.1.1 Problem description and formulation
-This first problem is simple and intuitive, with the solution known a priori, to demonstrate the simulation-based technique and validate all of our algorithms.
-In general, it has been suggested that the heating process should be done rapidly to avoid potential damages to the cells and maintain high viability (Terry et al., 2010; Baboo et al., 2019; Hunt, 2019; Uhrig et al., 2022), which is equivalent to the optimal control problem
+This first problem is simple and intuitive, with the solution known a priori, to demonstrate the simulation-based technique and validate all algorithms.
+In general, it has been suggested that the heating process should be done rapidly to avoid potential damage to the cells and maintain high viability (Terry et al., 2010; Baboo et al., 2019; Hunt, 2019; Uhrig et al., 2022), which is equivalent to the optimal control problem
 $\min_{\Theta_{b}(\tau)}\quad\tau_{f}$ (15)
 […]
-The solution to this optimal control problem is intuitive.
-To minimize the thawing time, the heater temperature should be fixed at its upper bound to provide the maximum heating power throughout the process.
+The solution to this optimal control problem is straightforward.
+To minimize the thawing time, the heater temperature should be fixed at its upper bound to deliver the maximum heating power throughout the process.
 4.1.2 DAE Reformulation
@@
 As a result, the equivalent system of DAEs for (15) is
-$\Theta_{b}(\tau)=1$ (16)
+$\Theta_{b}(\tau)=1,$ (16)
 […]
@@
 Every solution method gives the same optimal solution (Fig.
-2); the heater temperature $\Theta_{b}$ is at its upper bound of 1 all the time, ensuring that the algorithms and implementation are correct.
-The wall-clock time required for the simulation-based technique (sim_DAE) is more than an order of magnitude lower than every optimization-based approach (Table 1).
+2); the heater temperature $\Theta_{b}$ is at its upper bound throughout the process, indicating that our algorithms and implementations are correct.
+The wall-clock time required for the simulation-based technique (sim_DAE) is several orders of magnitude lower than those of the optimization-based approaches (Table 1).
 Optimization with GEKKO is the slowest method, followed by IPOPT, CasADi and fmincon.
 GEKKO’s ODE solvers cannot perform adaptive time-stepping, which could lead to significantly slower computation.
-Parallel computing fails to accelerate fmincon.
-The simulation-based approach is the most computationally efficient algorithm in this case study.
+Parallel computing fails to accelerate fmincon for this simple problem.
+Overall, the simulation-based approach is the most computationally efficient algorithm.
 Table 1: Computational performance of each solution method for Problem 1.
@@
 4.2.1 Problem description and formulation
-Besides using the fastest thawing as in Problem 1, some past studies considered cases in which the freezing and thawing rates are controlled, i.e., the rate of temperature change is kept constant (Seki and Mazur, 2008; Jang et al., 2017; Baboo et al., 2019; Bojic et al., 2021).
-As such, this problem focuses on controlling the rate of change in the average temperature during thawing by manipulating the heater temperature.
+Instead of using the maximum temperature as in Problem 1, some past studies considered cases in which the freezing and thawing rates are controlled, i.e., the rate of temperature change is kept constant (Seki and Mazur, 2008; Jang et al., 2017; Baboo et al., 2019; Bojic et al., 2021).
+As such, this problem focuses on controlling the rate of temperature change during thawing by manipulating the heater temperature.
 The definition of the average temperature is as given in (10).
-For a fixed heater temperature, the rate of temperature change is not constant (Figs.  2DE). Thus, the heater temperature needs to be manipulated. Consider the optimal control problem
+For a fixed heater temperature, the rate of temperature change is not constant (Figs. 2DE). Thus, the heater temperature needs to be manipulated. Consider the optimal control problem
 […]
 […]
-where $d\Theta_{\textrm{avg}}/d\tau$ is the rate of change in the average temperature and $(d\Theta_{\textrm{avg}}/d\tau)_{\mathrm{sp}}$ is the target value (setpoint) set to 0.04 for demonstration in this case study.
-For the optimization-based approaches, the control vector is parameterized using a piecewise linear function, with the number of control intervals $n_{c}=16$ (Appendix A).
+where $d\Theta_{\textrm{avg}}/d\tau$ is the rate of temperature change and $(d\Theta_{\textrm{avg}}/d\tau)_{\mathrm{sp}}$ is the target value (setpoint), set to 0.04 for demonstration in this problem.
+For the optimization-based approaches, the control vector is parameterized using a piecewise linear function, with the number of control intervals $n_{c}=16$ (Appendix B).
 4.2.2 DAE Reformulation
 For the simulation-based technique, first analyze the objective function in (17).
-This objective function is minimized when the rate of change in average temperature is equal to the setpoint.
+This objective function is minimized when the rate of temperature change is equal to the setpoint.
 Replacing the objective function in (17) with the algebraic equation $d\Theta_{\textrm{avg}}/d\tau=(d\Theta_{\textrm{avg}}/d\tau)_{\textrm{sp}}$ results in the system of DAEs
@@
 […]
-where $\Theta_{b}$ is now treated as a state. The differential index of (18) is 2, which can be solved using GEKKO as explained in Section 3.3.
+where $\Theta_{b}$ is now treated as a state.
+If the liquid domain used for calculating the average temperature includes the boundary point, the differential index of (18) is 1.
+If the boundary point is excluded, the index is 2.
+In any cases, (18) can be solved using GEKKO as explained in Section 3.3.
 4.2.3 Solution comparison
@@
 $||e||_{2}=\sqrt{\dfrac{1}{n_{k}}\sum_{k=1}^{n_{k}}\left(\left(\dfrac{d\Theta_{\textrm{avg}}}{d\tau}\right)_{\!k}-\left(\dfrac{d\Theta_{\textrm{avg}}}{d\tau}\right)_{\!\!\mathrm{sp}}\right)^{\!\!2}},$ (19)
-where $(d\Theta_{\textrm{avg}}/d\tau)_{k}$ is the rate of change in average temperature resulting from solving (17) or (18) evaluated at each time point $k$ in the time span [$\tau_{0},\tau_{f}$] and $n_{k}$ is the number of sampling time points.
-Small $||e||_{2}$ corresponds to the rate of change in average temperature being close to the target value, indicating that the heater temperature is optimal; i.e., the solution method is accurate.
+where $(d\Theta_{\textrm{avg}}/d\tau)_{k}$ is the rate of temperature change resulting from solving (17) or (18) evaluated at each time point $k$ in the time span [$\tau_{0},\tau_{f}$] and $n_{k}$ is the number of sampling time points.
+A small value of $||e||_{2}$ corresponds to the rate of temperature change being close to the target value, indicating that the heater temperature is optimal; i.e., the solution method is accurate.
 The optimal solutions to Problem 2 obtained from the optimization- and simulation-based approaches are illustrated in Fig.
 3.
-Most solution techniques predict the same heater temperature profile except the methods with fmincon that predict a slightly higher temperature (Fig.
+Most solution techniques predict the same heater temperature profile except the methods with fmincon, which predict a slightly higher temperature (Fig.
 3A).
 As time progresses, the heater temperature increases to compensate for a reduction in the temperature difference (driving force) between the heater and product.
-The rate of change in average temperature can be controlled nearly perfectly at 0.04 for IPOPT, CasADi, GEKKO, and the simulation-based method, while this value fluctuates up to about 0.046 for fmincon, implying that fmincon fails to converge to the optimal solution in this case (Fig.
+The rate of temperature change can be controlled nearly perfectly at 0.04 for IPOPT, CasADi, GEKKO, and the simulation-based method, while this value fluctuates up to about 0.046 for fmincon, implying that fmincon fails to converge to the optimal solution in this case (Fig.
 3B).
@@
 3C).
-Figure 3: Trajectories for the optimal (A) heater temperature, (B) rate of change in average temperature, and (C) average temperature simulated via the optimization- and simulation-based approaches in Problem 2.
+Figure 3: Trajectories of the optimal (A) heater temperature, (B) rate of temperature change, and (C) average temperature simulated via the optimization- and simulation-based approaches for Problem 2.
 Table 2: Computational performance and accuracy of each solution method for Problem 2.
@@
 For example, with $n=20$, there are 41 states for each control interval and 656 states for $n_{c}=16$.
-This large nonlinear optimization is computationally expensive to solve compared to other approaches.
+This large nonlinear optimization is computationally expensive to solve.
 In terms of accuracy, fmincon produces the largest error, which is about 30–40 fold higher than for the other techniques.
@@
 3.
-Although both fmincon and IPOPT employ the interior point methods, the detailed implementations (e.g., gradient/hessian calculation, scaling) are not exactly the same, which could result in different accuracy and performance (see detailed discussion and comparison of these solvers in Rojas-Labanda and Stolpe (2015)).
+Although both fmincon and IPOPT employ the interior point methods, the detailed implementations (e.g., gradient/hessian calculation, scaling) are not identical, which could result in different accuracy and performance (see detailed discussion and comparison of these solvers in Rojas-Labanda and Stolpe (2015)).
 The simulation-based and other optimization-based methods except fmincon have similar accuracy.
-Despite an increase in the differential index, the simulation-based approach is still much faster than the optimization-based techniques while giving the same level of accuracy.
+In this problem, the solution is now known a priori as in Problem 1, but the simulation-based approach is still much faster than the optimization-based techniques while giving the same level of accuracy.
 4.3 Problem 3: Control of the interface velocity
@@
 One of the primary interests in a Stefan problem is the moving solid-liquid interface.
-The literature has studied and shown the importance of interface position tracking and optimization in various applications (Hill and Wu, 1994; Brezina et al., 2018; Srisuma et al., 2023b, for example).
-This problem focuses on the control of the interface velocity, i.e., the evolution of a melting/freezing process.
+The literature has studied and shown the importance of interface tracking and optimization in various applications (Hill and Wu, 1994; Brezina et al., 2018; Srisuma et al., 2023b, for example).
+This problem focuses on controlling the interface velocity, i.e., the evolution of a melting/freezing process.
 Consider the optimal control problem
@@
 […]
-where $dS/d\tau$ is the interface velocity and $\left(dS/d\tau\right)_{\mathrm{sp}}$ is the target velocity set to $-0.1$ for demonstration in this case study.
-In Problem 1 (Fig.
-2BC), the velocity of the moving interface is not constant if the heater temperature is fixed.
-In this case study, the aim is to control the velocity of the moving interface to be constant at $(dS/d\tau)_{\mathrm{sp}}$ by manipulating the heater temperature throughout the process as formulated in (20).
-For the optimization-based approaches, the control vector is parameterized using a piecewise linear function, with the number of control intervals $n_{c}=12$ (see Appendix A).
+where $dS/d\tau$ is the interface velocity and $\left(dS/d\tau\right)_{\mathrm{sp}}$ is the target velocity, set to $-0.1$ for demonstration in this problem.
+In Problem 1 (Figs.
+2BC), the interface velocity is not constant when the heater temperature is fixed.
+In this problem, the aim is to control the interface velocity to be constant at $(dS/d\tau)_{\mathrm{sp}}$ by manipulating the heater temperature throughout the process as formulated in (20).
+For the optimization-based approaches, the control vector is parameterized using a piecewise linear function, with the number of control intervals $n_{c}=12$ (see Appendix B).
 4.3.2 DAE Reformulation
-To apply the simulation-based technique, a similar approach used in Problem 2 can be used, but here we instead enforce the moving interface velocity by modifying the interface equation (6) to be equal to $(dS/d\tau)_{\textrm{sp}}$.
+To apply the simulation-based technique, a similar approach used in Problem 2 can be used, but here we instead enforce the interface velocity by modifying the interface equation (6) to be equal to $(dS/d\tau)_{\textrm{sp}}$.
 Consequently, (20) is reformulated as the DAE system
@@
 The differential index of (21) is $n$, which is the number of grid points resulting from spatial discretization.
-Therefore, if the spatial discretization of the domain is made finer to increase numerical accuracy, the differential index of (21) increases.
-In this case, $n=20$, and hence (21) becomes an index-20 DAE system, making the problem even more complicated and difficult to solve compared to the index-2 system in Problem 2.
+Therefore, if the spatial discretization of the domain is made finer to increase numerical accuracy, the differential index of (21) also increases.
+In this case, $n=20$, and hence (21) becomes an index-20 DAE system, making the problem even more complicated and difficult to solve than Problems 1 and 2.
 To explore the possibility of using the index reduction technique, MATLAB’s reduceDAEIndex was used to reduce the differential index of (21).
 However, the index reduction process was not complete even after 20 hours of simulation.
-Obviously, this technique is not feasible for our problem as mentioned before in Section 3.3.
+Obviously, this technique is not feasible for our problem, as explained before in Section 3.3.
 Therefore, GEKKO’s DAE solver is needed.
@@
 4.
-The optimal heater temperature predicted by each technique is nearly identical except that there is a large drop around the end of the process for the fmincon cases (Fig.
+The optimal heater temperature predicted by each technique is nearly identical except that there is a large drop near the end of the process for fmincon (Fig.
 4A).
-The interface velocity can be controlled at about $-0.1$ all the time for IPOPT, CasADi, GEKKO, and the simulation-based technique, while some fluctuation is observed for fmincon (Fig.
+The interface velocity can be controlled at about $-0.1$ throughout the process for IPOPT, CasADi, GEKKO, and the simulation-based technique, while some fluctuation is observed for fmincon (Fig.
 4B).
@@
 4C).
-Figure 4: Trajectories for the optimal (A) heater temperature, (B) interface velocity, and (C) interface position via the optimization- and simulation-based approaches in Problem 3.
-For the computational performance, the simulation-based technique is about 60$\times$ to 900$\times$ faster than all of the optimization-based approaches (Table 3).
+Figure 4: Trajectories of the optimal (A) heater temperature, (B) interface velocity, and (C) interface position via the optimization- and simulation-based approaches for Problem 3.
+For the computational performance, the simulation-based technique is about 60$\times$ to 900$\times$ faster than all of the optimization-based approaches (Table 3), despite an increase in the differential index.
 IPOPT, GEKKO, and CasADi are comparable for the direct single shooting method, while the multiple shooting method is much slower.
 Parallel computing can significantly accelerate fmincon but is still much slower than the simulation-based approach.
-fmincon is the least accurate method; the error is higher than for the other approaches by more than an order of magnitude.
-Except for fmincon, the simulation-based approach has similar accuracy as the optimization-based approaches, with the error of about $2$$\times$$10^{-5}$.
+In terms of accuracy, the error of the simulation-based approach is about $2$$\times$$10^{-5}$, which is similar to those of the optimization-based approaches except for fmincon.
+With a relatively large error, fmincon is the least accurate method in this problem.
 Table 3: Computational performance and accuracy of each solution method for Problem 3.
@@
 The simulation-based approach, however, has one limitation associated with the DAE reformulation.
-As described in Section 3.3, the DAE reformulation is guaranteed to be optimal when at least one of the constraints, bounds, or algebraic equations resulting from reformulating the objective function is active.
+As described in Section 3.3, the DAE reformulation is guaranteed to be optimal when at least one of the inequality constraints, bounds, or algebraic equations resulting from reformulating the objective function is active.
 This is typically the case when the objective function of the original optimal control problem can be minimized by solving each subproblem resulting from control vector parameterization individually – in other words, local minimization at every control/time interval (subproblem) is equivalent to global minimization of the whole process (original problem).
@@
 4.4 Problem 4: Sensitivity analysis
-In Problems 2 and 3, the simulation-based approach is tested with one value of $(d\Theta_{\textrm{avg}}/d\tau)_{\mathrm{sp}}$ and $(dS/d\tau)_{\textrm{sp}}$, which are 0.04 and $-0.1$, respectively.
+In Problems 2 and 3, the simulation-based approach is tested with one set of values of $(d\Theta_{\textrm{avg}}/d\tau)_{\mathrm{sp}}$ and $(dS/d\tau)_{\textrm{sp}}$, which are 0.04 and $-0.1$, respectively.
 To demonstrate the robustness of our framework, this section conducts a sensitivity analysis by varying $(d\Theta_{\textrm{avg}}/d\tau)_{\mathrm{sp}}$ and $(dS/d\tau)_{\textrm{sp}}$ and employs the simulation-based approach to solve the problems.
@@
 5 and 6).
-The rate of change in average temperature and interface velocity are at the setpoints except for $(dS/d\tau)_{\textrm{sp}}=-0.15$.
+The rate of temperature change and interface velocity are at the setpoints except for $(dS/d\tau)_{\textrm{sp}}=-0.15$.
 An interface velocity of $-0.15$ is too fast for the given heater temperature, so the upper bound is active, and that the heater temperature cannot be increased further to achieve the target velocity.
-The target velocity that is larger (smaller in magnitude) than the peak of Fig.
+A target velocity that is larger (smaller in magnitude) than the peak of Fig.
 2C, which about $-0.12$, will never violate the bound (see Appendix C for a formal proof).
-For cases where the bound is active, a simulation switching technique can be used to transition between active constraints, resulting in a hybrid discrete/continuous dynamic simulation (Feehery and Barton, 1998; Barton et al., 2000; Berliner et al., 2022).
+For cases where the bound is active, a simulation-switching technique can be used to transition between active constraints, resulting in a hybrid discrete/continuous dynamic simulation (Feehery and Barton, 1998; Barton et al., 2000; Berliner et al., 2022).
 First, initialize a DAE simulation as usual.
@@
 (2023a)).
-GEKKO’s DAE solvers, on the other hand, do not have a built-in option for handling switches, and so this process needs to be executed manually (see our software/code for implementation).
-This switching technique can also be used for path constraints.
-Figure 5: Trajectories for the optimal (A) heater temperature, (B) rate of change in average temperature, and (C) average temperature simulated via the simulation-based approach at four different values of $(d\Theta_{\textrm{avg}}/d\tau)_{\mathrm{sp}}$ in Problem 4.
-Figure 6: Trajectories for the optimal (A) heater temperature, (B) interface velocity, and (C) interface position simulated via the simulation-based approach at four different values of $(dS/d\tau)_{\textrm{sp}}$ in Problem 4.
+GEKKO’s DAE solvers, on the other hand, do not have a built-in option for handling switches, and so this process needs to be executed manually.
+This switching technique can also be extended to handle path constraints.
+Figure 5: Trajectories of the optimal (A) heater temperature, (B) rate of temperature change, and (C) average temperature simulated via the simulation-based approach at four different values of $(d\Theta_{\textrm{avg}}/d\tau)_{\mathrm{sp}}$ for Problem 4.
+Figure 6: Trajectories of the optimal (A) heater temperature, (B) interface velocity, and (C) interface position simulated via the simulation-based approach at four different values of $(dS/d\tau)_{\textrm{sp}}$ for Problem 4.
 With this capability, the benefit of the simulation-based approach is even more substantial in applications where a large number of simulation runs are required, e.g., parametric studies and design optimization.
@@
 The optimal control and state vectors are obtained via simulation of the resulting DAE system with the selected DAE solver, eliminating the need for an optimization solver and thus greatly accelerating the computation.
-Our proposed framework was demonstrated and benchmarked against a variety of optimization-based approaches – IPOPT, fmincon, CasADi, and GEKKO – for case studies involving the optimal control of a Stefan problem for cell thawing.
-The case studies include the minimization of the thawing time, the control of the average temperature, and the control of the interface velocity, transformed into DAE systems of index 1, 2, and 20, respectively.
-At the same level of accuracy, the simulation-based approach is more than an order of magnitude faster than every optimization-based method in all cases.
-The simulation-based approach offers an efficient and reliable framework for solving the optimal control problems, serving as a promising alternative to the traditional techniques in applications where speed is crucial, e.g., real-time online model predictive control.
-Some future directions of interest are to (1) generalize the approach and (2) improve the computational performance.
-The ultimate goal is a highly efficient optimal control solver applicable to problems with a variety of objective functions and constraints that arise in multiple fields.
-Implementation would be simplified if high-index DAE solvers become available that have built-in options for mixed continuous-discrete simulation.
-Data Availability
-All software and data will be made available upon publication of the manuscript.
+Our proposed framework, the simulation-based approach, is demonstrated and benchmarked against a variety of optimization-based approaches – IPOPT, fmincon, CasADi, and GEKKO – for case studies involving the optimal control of a Stefan problem for cell thawing.
+The case studies include the minimization of the thawing time, the control of the average temperature, and the control of the interface velocity, covering both index-1 and high-index DAE systems.
+The simulation-based approach is faster than every optimization-based method by more than an order of magnitude while giving similar/better accuracy in all cases.
+The current simulation-based approach, however, is limited to problems where at least one of the inequality constraints, bounds, or algebraic equations resulting from reformulating the objective function is active.
+Some future directions of interest are to (1) generalize the approach for more complicated problems, objective functions, and constraints that arise in multiple fields and (2) improve the computational performance with more efficient solvers and implementation.
 Acknowledgments
@@
 Appendices
-Appendix A Control Vector Parameterization
-A.1 Piecewise constant and linear controls
+Appendix A Solver Options
+There are a large number of options for optimization, ODE, and DAE solvers, which could lead to different accuracy and computational performance.
+This section describes and justifies the choice of important solver options used in our simulations.
+The default values are used for options that are not listed here.
+For all case studies, the selected DAE solvers are MATLAB’s ode15s for index-1 DAEs and GEKKO’s DAE solver (Python) for high-index systems.
+The differential index of DAEs can be checked using MATLAB’s reduceDAEIndex.
+The wall-clock times (aka wall times, clock times) can be measured using the tic and toc functions, with each simulation repeated at least 10 times until the standard deviation of the measured wall times is smaller than 5% for consistent results.
+The initial guess of $\Theta_{b}$ is set to 0.5, the average dimensionless temperature, so that every solution method is initialized and compared on the same basis.
+For the optimization-based approaches, the optimality tolerance is set to $10^{-7}$ for all optimization solvers, including IPOPT, fmincon, CasADi, and GEKKO.
+The integration tolerance for ode15s is $10^{-10}$.
+Using higher values for the optimality and integration tolerances could lead to inaccurate solutions, while using tighter values does not improve the accuracy significantly.
+For the objective functions (17) and (20) in Problems 2 and 3, the derivatives are approximated using a finite difference scheme, whereas the integrals can be approximated using a Riemann sum.
+The time step $\Delta\tau$ is set to 0.05.
+While ode15s uses adaptive time-stepping, this option is not available in GEKKO; i.e., users manually specify a vector of time points for ODE integration.
+Hence, there is no ideal comparison between GEKKO’s ODE solvers and MATLAB’s ode15s.
+The fixed time step $\Delta\tau=0.05$ is specified for ODE integration in GEKKO, with five collocation points (INODES = 5) for each time interval.
+This time step is chosen to be consistent with the value used for a finite difference and Riemann sum approximation mentioned in the previous paragraph, while the number of collocation points is selected such that that the temperature and interface position simulated by GEKKO’s ODE solver have the same level of accuracy as those simulated by ode15s.
+For the simulation-based approach, ode15s is used for an index-1 system (Problem 1), with the exact same solver options as mentioned above.
+GEKKO’s DAE solver is used for high-index systems (Problems 2 and 3).
+As a reformulated DAE system is different from the original ODE system, the time step needs to be adjusted accordingly.
+The time step is selected such that the accuracy of the simulation-based solution measured by $||e||_{2}$ is on the same order of magnitude as that of the optimization-based solution, giving $\Delta\tau=0.12$ and $\Delta\tau=0.37$ for Problems 2 and 3, respectively.
+The accuracy of the optimization- and simulation-based approaches is measured using $||e||_{2}$ defined by (19) and (22).
+To calculate $||e||_{2}$, the derivatives are approximated using a finite difference scheme with $\Delta\tau=0.05$.
+Appendix B Control Vector Parameterization
+This appendix describes the choice of control vector parameterization used for the optimization-based approaches.
+B.1 Piecewise constant and linear controls
 Piecewise constant and linear controls are most common in optimal control (Nolasco et al., 2021).
 This section justifies the use of piecewise linear controls in Problems 2 and 3 for the optimization-based approaches, in comparison to piecewise constant controls.
-This comparison considers the most complex problem (Problem 3), with opt_Ipopt as the solver.
-Table 4: Comparison between the errors $||e||_{2}$ resulting from piecewise constant and linear control vector parameterization for different values of control intervals $n_{c}$.
+This comparison considers the most complex problem, Problem 3, and uses opt_Ipopt as the solver.
+Table 4: Comparison between the errors $||e||_{2}$ of piecewise constant and linear control vector parameterization for different values of control intervals $n_{c}$.
 $n_{c}$ Error measured by $||e||_{2}$
@@
 As a result, a piecewise linear control is selected in this work.
-A.2 Number of control intervals
+B.2 Number of control intervals
 The accuracy of the optimization-based solution is influenced by the number of control intervals $n_{c}$.
 This section justifies those numbers by investigating the effect of $n_{c}$ on the accuracy of the solution.
-The optimal control problems defined by (17) and (20) were solved via opt_Ipopt and $||e||_{2}$ was calculated for different values of $n_{c}$.
-Logically, the accuracy of the solution should improve, i.e., smaller $||e||_{2}$, when $n_{c}$ increases due to finer discretization.
+The optimal control problems defined by (17) and (20) are solved via opt_Ipopt and $||e||_{2}$ is calculated for different values of $n_{c}$.
+Logically, the accuracy of the solution should improve (i.e., $||e||_{2}$ decreases) when $n_{c}$ increases, due to finer discretization.
 From Fig.
 7, the accuracy of the solution does not change significantly after $n_{c}$ reaches some certain value, and so this threshold should be chosen for control vector parameterization.
-As a result, $n_{c}$ was set to 16 for Problem 2 (Fig.
+As a result, $n_{c}$ is set to 16 for Problem 2 (Fig.
 7A) and 12 for Problem 3 (Fig.
@@
 Figure 7: Errors measured by $||e||_{2}$ for different values of control intervals $n_{c}$ for (A) Problem 2 and (B) Problem 3.
-Appendix B Solver Options
-There are a large number of options for optimization, ODE, and DAE solvers, which could lead to different accuracy and computational performance.
-This section describes and justifies the choice of important solver options used in this work.
-The default values were used for the options not mentioned here.
-For all case studies, index-1 DAEs were solved by ode15s in MATLAB, whereas high-index DAEs were handled by GEKKO in Python.
-The differential index of DAEs was checked using reduceDAEIndex.
-The wall-clock times (aka wall times, clock times) were measured using the tic and toc functions, with each simulation repeated at least 10 times for consistent results, as indicated by the standard deviation of the measured wall times being smaller than 5%.
-For fair comparison of computational performance, the initial guess of $\Theta_{b}$ was set to the average value of 0.5 in all cases.
-For the optimization-based approaches, the optimality tolerance was set to $10^{-7}$ for all optimization solvers, including IPOPT, fmincon, CasADi, and GEKKO.
-The integration tolerance for ode15s was set to $10^{-10}$.
-Using higher values for the optimality and integration tolerances could lead to inaccurate solutions, while using tighter values does not improve the accuracy significantly.
-For the objective functions (17) and (20) in Problems 2 and 3, the derivatives were approximated using a finite difference scheme, whereas the integrals can be approximated using a Riemann sum.
-The time step $\Delta\tau$ was set to 0.05.
-While ode15s uses adaptive time-stepping, this option is not available in GEKKO; i.e., users have to specify a vector of time points for ODE integration.
-Hence, there is no ideal comparison between GEKKO’s ODE solvers and MATLAB’s ode15s.
-The fixed time step $\Delta\tau=0.05$ was specified for ODE integration in GEKKO, with five collocation points (INODES = 5) for each time interval.
-This time step was chosen to be consistent with the value used for a finite difference and Riemann sum approximation mentioned in the previous paragraph, while the number of collocation points was selected such that that the temperature and interface position simulated by GEKKO’s ODE solver have the same level of accuracy as those simulated by ode15s.
-For the simulation-based approach, ode15s was used for an index-1 system (Problem 1), with the exact same solver options as mentioned above.
-GEKKO’s DAE solver was used for high-index systems (Problems 2 and 3).
-As a reformulated DAE system is different from the original ODE system, the time step needs to be adjusted accordingly.
-The time step was selected such that the accuracy of the simulation-based solution measured by $||e||_{2}$ is on the same order of magnitude as that of the optimization-based solution, giving $\Delta\tau=0.12$ and $\Delta\tau=0.37$ for Problems 2 and 3, respectively.
-The accuracy of both the optimization- and simulation-based approaches is measured using $||e||_{2}$ defined by (19) and (22).
-To calculate $||e||_{2}$, the derivatives were approximated using a finite difference scheme with $\Delta\tau=0.05$.
-Appendix C Bounds on the Interface Velocity and Rate of Change in Temperature
+Appendix C Bounds on the Interface Velocity and Rate of Temperature Change
 We show in Fig.
@@
 A formal proof is given in this section.
-To obtain a closed-form solution for the proof, the heat conduction in the liquid domain is assumed to be quasi-steady.
-This approximation is accurate for problems associated with phase change as most of the heat transfer occurs at the moving interface.
+To obtain a closed-form solution for the proof, heat conduction in the liquid domain is assumed to be quasi-steady.
+This approximation is accurate for problems associated with phase change in general as the sensible heat is much smaller than the latent heat.
 In this case, the closed-form solution for the interface velocity is
 $\frac{dS}{d\tau}=-\frac{k_{2}U\Theta_{b}(T_{0}-T_{m})}{\rho L\alpha_{1}S\left(U\ln(1/S)+k/b\right)},$ (23)
-where all the parameter description and values can be found in Srisuma et al.
+where the parameter values and description can be found in Srisuma et al.
 (2023b).
-From (23), the interface velocity $(dS/d\tau)$ is a monotonic function of the heater temperature $(\Theta_{b})$.
-Hence, the condition in which the bounds on the heater temperature are not violated can be obtained by substituting the upper bound $\Theta_{b}=1$ into (23), which is
-$\left(\dfrac{dS}{d\tau}\right)_{\!\!\mathrm{sp}}>-\frac{k_{2}U(T_{0}-T_{m})}{\rho\alpha_{1}\Delta H_{\!f}S\left(U\ln(1/S)+k/b\right)}.$ (24)
+From (23), the interface velocity $dS/d\tau$ is a monotonic function of the heater temperature $\Theta_{b}$.
+Substituting the upper bound $\Theta_{b}=1$ into (23) yields
+$\left(\dfrac{dS}{d\tau}\right)_{\!\!\mathrm{sp}}>-\frac{k_{2}U(T_{0}-T_{m})}{\rho\alpha_{1}\Delta H_{\!f}S\left(U\ln(1/S)+k/b\right)},$ (24)
+which gives the limit on the target velocity that can be specified without violating the upper bound on the heater temperature.
 With the parameter values in Srisuma et al.
@@
 This conclusion agrees with the results presented in Problems 3 and 4, in which the bounds are not violated when $(dS/d\tau)_{\mathrm{sp}}$ is higher than $-0.13$, whereas the upper bound is active when $(dS/d\tau)_{\mathrm{sp}}$ is $-0.15$.
-A similar strategy can be used for analyzing the temperature change, e.g., using a thermal lumped capacity approximation.
+A similar strategy can be used to analyze the temperature change, e.g., using a thermal lumped capacity approximation.
 Note that our simulation-based approach does not require to know a priori if the bounds will be active or not, as the upper and lower bounds can be handled as shown in Problem 4.
+Data Availability
+Software and data used in this work are available at https://github.com/PrakitrSrisuma/simDAE-optimalcontrol.